Two angles are complementary. The sum of the measure of the first angle and half the second angle is 89.5 degrees. What is the measure of the smaller angle? What is the measure of the other angle?

Please help. Thank you

Found 2 solutions by shree840, htmentor:
Answer by shree840(260)

complementary means two angle combined makes 90 degrees
first and half the second angle are 89.5
without going into details 89 and 1 is only possible.
These are not angles of a triangle i hope.

Answer by htmentor(788)

Let x = the first angle
Then the second angle = 90-x, since the two angles are complementary
We can write the following equation:
x + (90-x)/2 = 89.5
x/2 + 45 = 89.5
x = 89
The other angle = 90 - x = 90 - 89 = 1
So, the smaller angle = 1 deg, the larger angle = 89 deg